Find the equation of the parabola.
Focus (5,-2), directrix x=-1

Answer by lwsshak3(11628)

Find the equation of the parabola.
Focus (5,-2), directrix x=-1
This is an equation of a parabola that opens rightwards.
Its standard form: (y-k)^2=4p(x-h), (h,x)=(x,y) coordinates of the vertex
For given parabola:
y- coordinate of vertex=-2
x-coordinate of vertex=2 (midpoint of -1 and 5)
vertex:(2,-2)
axis of symmetry: y=-2
p=3 (distance from focus or directrix to vertex on the axis of symmetry)
equation:
(y+2)^2=12(x-2)